OpenStudy (anonymous):

Can someone help me solve (-32) - 29

OpenStudy (anonymous):

-32 - 29 = -3

OpenStudy (anonymous):

-61

OpenStudy (cwrw238):

imagine descending 32 steps and then another 29 - how many '-' s is that?

OpenStudy (anonymous):

3

OpenStudy (cwrw238):

or its -32 degrees in deepest Siberia and night falls and temperature goes down another 29 degrees!!

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-3 :)

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

k whats 32 - 29?

OpenStudy (anonymous):

i dont think it is 3

OpenStudy (anonymous):

-61

OpenStudy (anonymous):

@ ktnguyen1 is correct

OpenStudy (anonymous):

@ktnguyen1 how did ya get it

OpenStudy (anonymous):

now the greater value is -32 so the greater value has to have the same exponenent as the answer

OpenStudy (anonymous):

so -32 - 29 is -3

OpenStudy (anonymous):

let me get this straight. two negative is a positive right

OpenStudy (anonymous):

both are not negatives

OpenStudy (anonymous):

one is -32 and the other is +29

OpenStudy (anonymous):

that means we add the two numbers but keep the bigger number's sign

OpenStudy (anonymous):

the - prabuably got you confused

OpenStudy (anonymous):

yes excacly

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

+ x - = -

OpenStudy (cwrw238):

2 negatives make a positive ehen multiplying eg -2 * -2 = +4 or in the a case like 2 - (-4) = 2+ 4 = 6 but -4 -6 = -10 or (-4)-6 = -10

OpenStudy (anonymous):



OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

so isn it -61 ?